OpenStudy (anonymous):

Two trees stand side-by-side. The shorter tree is feet tall. Its shadow is 9 feet long. The taller tree is feet tall. How long is its shadow?

OpenStudy (anonymous):

Missing a little data in the problem....

OpenStudy (anonymous):

Let the shadow be x and use trigonometry and the tangent ratio: 18/x = 4/9 x = 40.5 feet